Declining perpetuities and annuities. You own an oil pipeline that will generate a $2 million cash return over the coming year. The pipeline’s operating costs are negligible, and it is expected to last for a very long time. Unfortunately, the volume of oil shipped is declining, and cash flows are expected to decline by 4% per year. The discount rate is 10%.

  1. What is the PV of the pipeline’s cash flows if its cash flows are assumed to last forever?

  2. What is the PV of the cash flows if the pipeline is scrapped after 20 years?
Answer #1

1.

Present valuenof perpetuity=Cas flow next year/(discount rate-growth rate)=2/(10%-(-4%))=14.29

2.

Present value of annuity with growth rate g=Cash flow next year/(1+discount rate)*(1-((1+g)/(1+discount rate))^n)/(1-(1+g)/(1+discount rate))=2/1.1*(1-(0.96/1.1)^20)/(1-0.96/1.1)=13.35